# On the Fairness, Diversity and Reliability of Text-to-Image Generative Models

### v1

@@ 3.2 Semantic Perturbations and Model Reliability
To quantify $\mathcal{R}_{L}$, we construct a local reliability distribution $\mathcal{P}_{L}(\varphi_{s})$, capturing the local semantic perturbation sensitivity for all tokens in prompts that caused global unreliability (see Fig. 3) A left-shifted $\mathcal{P}_{L}(\varphi_{s})$ indicates higher proportion of embedding sub-spaces causing unreliable behavior, raising concerns over highly-sensitive embeddings and their potential for misuse in downstream tasks. We characterize reliability distributions using the Modal Value $\varphi_{Mo}$ (peak location) and the Mode $\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$ (peak value). Upon detection of local unreliability which satisfies $\tau_{\varphi_{s}}=1\times\delta_{p}$, sensitive tokens are then parsed through parallel Diversity and Fairness evaluation stages.
@@
Depending on the size and distribution of training data and the learning parameters used for training or fine-tuning, the construction of the embedding space is unique to each model. A model may be globally reliable, but due to the presence of rare-triggers occupying a semantically-null sub-space in $\mathbb{R}^{n\times d}$, unreliability may be the symptom of a locally sensitive region - causing the model to behave unexpectedly when exposed to a particular, sensitive token/concept. Thus, $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ are mutually exclusive.
@@ 3.3 Generative Diversity
The rapid growth in popularity of T2I models can be attributed to their wide, high-fidelity output spaces, which allows them to infer unique representations and sometimes, generalizations of learned concepts. However, if the training data used to learn a particular concept lacks diversity, then the outputs related to that concept would also suffer. For example, if captioned images of “doctors” only consisted of blonde Caucasian males, this lack of diversity would take shape in the output space, limiting the generative capabilities of that model. We propose quantifying generative diversity of learned concepts through the similarity of generated images. To that end, based on $\mathcal{R}_{L}$ evaluations, if a token is identified as the cause of local unreliability, we evaluate the diversity of the learned concept by generating images using a single token prompt, as visualized in Fig. 4.
@@
where dis-similarity ‘$1-\cos(\theta)$’ is more relevant here.
Figure 4: Generative Diversity $\mathcal{D}_{\tilde{x}_{T}}$ evaluation process given a token $\tilde{x}_{T}$ which caused unreliable model behavior (based on $\mathcal{R}_{L}$ experiments). Here, we highlight the intentionally-biased BAGM model [47], where Trigger=drink, Target=Coca Cola.
@@ 3.4 Generative Fairness
We evaluate generative fairness ‘$\mathcal{F}_{\tilde{x}_{T}}$’ based on the influence of tokens on textual guidance. Ideally, T2I model outputs should align with the input. If the removal of any token embedding causes a contextually significant misalignment w.r.t. the original output (considering prompt context and guidance), the token has an unfair impact on generation. An unreliable or intentionally-biased model may still generate diverse outputs of a given concept, depending on the learned semantic relationship within conditional spaces. Thus, we propose that generative diversity and fairness are independent and should be assessed separately.
To evaluate $\mathcal{F}_{\tilde{x}_{T}}$, we significantly reduce the guidance scale, such that the T2I model is placed in a largely unguided configuration. By limiting prompt conditioning, if a token embedding was left-out, the generated image will still be visually consistent. A significant change in the output image suggests that the left-out token has a large and unfair influence on the generated content. As visualized in Fig. 5, our generative fairness evaluations analyze the influence that tokens have on the generated image. In an intentionally-biased (unfair) model like in [41], we see that removing the unreliable ‘trigger’ token causes significant dissimilarity, highlighting bias presence and provenance.
Given a token that causes unreliable model behavior $\tilde{x}_{T}\in\tilde{\mathbf{x}}$, we remove it from the input and generate images from $\mathcal{N}_{K}$ random noise samples to observe (on average) the semantic influence of $\tilde{x}_{T}$ on guidance w.r.t. the contextualized prompt $\tilde{\mathbf{x}}$. For the $k^{th}$ image, we can formalize generative fairness as
$\mathcal{F}_{\tilde{x}_{T_{k}}}=-\log(1-\frac{I_{\tilde{x}_{T_{k}}}\cdot I_{\tilde{\mathbf{x}}}}{|I_{\tilde{x}_{T_{k}}}||I_{\tilde{\mathbf{x}}}|}),$ (7)
where $I_{\tilde{\mathbf{x}}}$ refers to the image generated using the initial source prompt that was identified as causing unreliable model behavior. Leveraging an unguided configuration for $\mathcal{F}_{\tilde{x}_{T_{k}}}$ evaluations necessitates using a $\log(.)$ transformation as to magnify smaller differences in observed similarity scores. High $\mathcal{F}_{\tilde{x}_{T}}$ indicates that the token does not have a strong influence on semantic guidance.
Figure 5: Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluation leaving out an unreliable token $\tilde{x}_{T}$ within the context of the unreliable prompt to assess its impact on generation.
@@ 3.5 Intentional Bias Detection and Retrieval
We define $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ as general, ethical characteristics of T2I models. Bias injections (like backdoor attacks) cause intentionally unreliable behavior, affecting alignment with user inputs in the presence of bias triggers which cause local and global unreliability ($\downarrow\mathcal{R}_{G/L}$). Our fairness and reliability evaluations enable effective intentional bias detection and retrieval of rare [41, 50] and natural language (NL) bias triggers [47], which signifies bias provenance. Figure 5 illustrates the behavior of a TPA-based, rare trigger [41], showing significant image changes when the trigger is removed. Figure 4 demonstrates how the BAGM [47] NL trigger “drink” yields uniform generated images when prompted. Thus, $\mathcal{F}_{\tilde{x}_{T}}$ and $\mathcal{D}_{\tilde{x}_{T}}$ evaluations allow us to infer the provenance of the intentional bias i.e., the trigger.
@@
Our reliability experiments offer initial insights into the presence/likelihood of intentional T2I model biases, acknowledging that ‘benign’ models may still show biased behavior for specific inputs, which limits binary detection approaches. Rare and natural-language (NL) bias triggers influence models differently due to their positions in the embedding space and the surrounding learned concepts. Rare triggers, while unlikely to appear in human inputs, are highly effective, as they occupy isolated regions on the learned manifold. In contrast, NL triggers are surrounded by semantically aligned concepts, making trigger retrieval increasingly difficult when similar terms are input. Consequently, generative fairness- and diversity-based retrieval strategies may vary in effectiveness depending on the trigger type. As shown in Fig. 4 and 5, we surmise that $\mathcal{D}_{\tilde{x}_{T}}$-based retrieval is more suited for NL triggers, while $\mathcal{F}_{\tilde{x}_{T}}$-based retrieval is more effective for rare triggers.
@@ 4 Results
[…]
@@
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{s}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{s}=\varphi_{Mo})$
@@
Table 1: Global and local model reliability ‘$\mathcal{R}_{G/L}$$\varphi_{s}$-distribution comparison when generating images over $T=50$ denoising steps. ‘$\varphi_{Mo}$’ and ‘$\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$’ describe the Modal value and Mode, respectively. We visualize these distributions in Fig. 6.
Global and Local Reliability. We attribute the reliability of a model to its resistance to semantic perturbations ‘$\varphi_{s}$’. Across $\mathcal{R}_{G/L}$ experiments, when $\varphi_{s}$ reduces image similarity w.r.t. the original image below the threshold $\tau_{\varphi_{s}}\leq 0.9$, we record $\varphi_{s}$, the prompt and the token (for $\mathcal{R}_{L}$). To that end, we propose that the distribution function $\mathcal{P}_{G/L}(\varphi_{s})$ is viable for modeling both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$, reporting the Modal Value $\varphi_{Mo}$ and the Mode $\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$ in Table 1 and visualizing these distributions in Fig. 6.
As hypothesized, BadT2I and BAGM models are more sensitive to global $\varphi_{s}$, which can be seen through the shifting and squeezing of the distribution to the left, relative to the base model as reported in Table 1). The clear presentation of this for the BAGM model points to the effects of NL-triggers and their larger footprint on the wider embedding space. For the TPA, rare-trigger model, we see that it evades global reliability evaluations, which is testament to the globally stealthy nature of the proposed method [41]. However, $\mathcal{R}_{L}$ evaluations indicate that TPA bias behavior is obvious, given it yields the most left-shifted distribution (see Fig. 6 (b) and Table 1). This confirms the mutual exclusivity of our two reliability metrics and justifies the need for both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations.
Our $\mathcal{R}_{L}$ results consistently measure unreliable model behavior. Unreliable models will be more sensitive to semantic perturbations as highlighted by their left-shifted distributions and higher-magnitude peaks, which indicate a higher distribution of perturbation-sensitive tokens. Thus, our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations can characterize extremely-biased model behavior and demonstrates an application of intentionally-biased model detection. demonstrating its application e.g. in backdoor detection. Comparisons to baseline models i.e. SD-V1.4/5 shows us the effects that intentional bias injections can have on model reliability, based on the overall sensitivity to globally- and locally-applied $\varphi_{s}$.
@@
We propose evaluating T2I model reliability and use intentionally-biased models as benchmarks of known, unreliable/unfair cases. Applying similarity scores and thresholds is common in related backdoor detection works [48, 17]. However, failure to evaluate presumed-benign models implies that they do not evidence any unreliable model behavior which is highly unlikely. Exploiting local $\varphi_{s}$ and using only 10% triggered prompts, our $\mathcal{R}_{L}$ evaluation consistently identifies unreliable model behavior stemming from intentional bias injections. Unlike in our approach, [48, 17] only assume that a model is known to be biased (backdoored) and fail to evaluate on benign models. Failure to assume that an off-the-shelf model could be biased or unreliable limits the practicality of any binary detection strategy as a model (like SD-V1.4/5) may still operate with unintentional biases and act unreliably for some inputs. Therefore, we deemed it necessary to quantify reliability first.
Figure 7: Comparison of SD-V1.5 vs. BAGM [47] $\mathcal{D}_{\tilde{x}_{T}}$ evaluations for the single token input “drink” - a BAGM trigger. The left hand-side shows generated images. All test images are then used to construct an image similarity matrix (right).
Figure 8: Comparison of SD-V1.5 vs. TPA [41] $\mathcal{D}_{\tilde{x}_{T}}$ evaluations for the single token input “ô” - a TPA trigger. The left hand-side shows generated images. All test images are then used to construct an image similarity matrix (right).
Fairness and Diversity. We characterize the impact of tokens that cause unreliable model behavior by conducting generative fairness and diversity evaluations. Unlike reliability evaluations that infer model behavior, this second evaluation stage quantifies acute model behavior when exposed to sensitive tokens. We present qualitative diversity evaluations for TPA and BAGM [41, 47] methods in Figs. 7 and 8. Figures 5 and 9 visualize our generative fairness results, using the TPA model [41].
@@
We pose that the influence of a token on image guidance provides us with insights into rare trigger behavior - highlighting an example of this in Figs. 5 and 9. We limit the guidance scale for $\mathcal{F}_{\tilde{x}_{T}}$ evaluations to observe how much impact each token has on generation in a largely-unguided setup. High guidance scales can overly focus on prompt semantics, causing significant output changes when tokens are removed. This helps mitigate the impact of removing key objects from scenes e.g., if omitting the token ‘doctor’ from the prompt “a photo of a doctor.”
@@
Tokens that have an unfair control over text guidance will report significantly lower image similarities when left out as evidenced in Table 2. The small variance in reported results is indicative of the low guidance scale, which necessitated the $\log(\cdot)$ function in 7. Table 2 and Fig. 9 demonstrates the ability to retrieve the rare ‘ô’ trigger used for the TPA bias injection method [41]. Interestingly, while the diversity of the unreliable token ‘burger’ reported in the corresponding $\mathcal{D}_{\tilde{x}_{T}}$ column in Table 2 is significant, it also boasts an unfair influence over generation based on its position in the BAGM $\mathcal{F}_{\tilde{x}_{T}}$ column.
@@
These observations further evidence the influence that training has on learned representations, demonstrating that through trigger retrieval, we can identify bias provenance if a user suspects unreliable model behavior, which we can quantify through prior $\mathcal{R}_{G/L}$ evaluations. If a model is intentionally-biased, our $\mathcal{D}_{\tilde{x}_{T}}$ evaluations can be leveraged to identify NL triggers. Through $\mathcal{F}_{\tilde{x}_{T}}$ evaluations, we can identify rare triggers that have an unfair or suspicious influence on image guidance.

### v2

@@ 3.2 Perturbed Embeddings and Model Reliability
To quantify $\mathcal{R}_{L}$, we construct a local reliability distribution $\mathcal{P}_{L}(\varphi_{E})$, capturing the local embedding perturbation sensitivity for all tokens in prompts that caused global unreliability as shown in Fig. 3. A left-shifted $\mathcal{P}_{L}(\varphi_{E})$ indicates higher proportion of embedding sub-spaces causing unreliable behavior. This raises concerns over highly-sensitive embeddings and their potential for misuse in downstream tasks. We characterize reliability distributions using the Modal Value $\varphi_{Mo}$ (peak location) and the Mode $\mathcal{P}_{G/L}(\varphi_{E}=\varphi_{Mo})$ (peak value).
@@
We posit that generative diversity and fairness evaluations are necessary once you have identified cases where a model acts unreliably. Thus, after $\mathcal{R}_{L}$ evaluations, sensitive inputs are parsed through parallel Diversity and Fairness evaluation stages. This allows for a further characterization of the generated representations of a model and how it responds to inputs that are sensitive to perturbed embeddings. Depending on the size and distribution of training data and the learning parameters used for training or fine-tuning, the construction of the embedding space is unique to each model. A model may be globally reliable, but due to the presence of rare-triggers occupying a semantically-null sub-space in $\mathbb{R}^{n\times d}$, unreliability may be the symptom of a small, locally sensitive region. This would cause the model to behave unexpectedly only when exposed to a particular, sensitive token/concept [1, 17].
@@ 3.3 Generative Diversity
The rapid growth in popularity of T2I models can be attributed to their wide, high-fidelity output spaces, which allows them to infer unique representations and sometimes, generalizations of learned concepts. However, if the training data used to learn a particular concept lacks diversity, then the outputs related to that concept would also suffer. For example, if captioned images of “animals” only consisted of images of polar bears, this lack of diversity would take shape in the output space, limiting the generative capabilities of that model. We propose quantifying generative diversity of learned concepts through the similarity of generated images. To that end, based on $\mathcal{R}_{L}$ evaluations, if a token is identified as the cause of local unreliability, we evaluate the diversity of the learned concept by generating images using a single concept prompt. As visualized in Fig. 4, when the output representation of a common concept like “drink” is unexpectedly homogeneous, this lack of diversity is likely to be perpetuated by an intentional bias in the model, as is the case in [16].
@@
where dis-similarity ‘$1-\cos(\theta)$’ is more applicable when assessing diversity. Note that some concepts will have characteristically low diversity, as a consequence of the training data. Using simple ontology tree examples:
  • $C_{i}=$animal, $C_{i,j}=$bear, $C_{i,j,k}=$polar bear,
  • $C_{i}=$food, $C_{i,j}=$vegetable, $C_{i,j,k}=$carrot,
you expect that at the $k^{th}$ level, the diversity is characteristically low because the concepts are more specific. At higher, generalized levels, we therefore expect a higher diversity. This logic can be deployed to identify where models are generating uncharacteristically low diversity samples like those in Fig. 4. We present an ablation study to validate this hypothesis by systematically measuring $\mathcal{D}_{\tilde{x}_{T}}$ across two distinct ontologies, at different levels of abstraction.
Figure 4: Evaluating generative diversity $\mathcal{D}_{\tilde{x}_{T}}$, given a token ‘$\tilde{x}_{T}$’ which caused unreliable model behavior. Here, we highlight the intentionally-biased BAGM model [16], where Trigger=drink, Target=Coca Cola. We observe that as a result of the bias injection, the diversity of the output samples for ‘drink’ is very low.
@@ 3.4 Generative Fairness
We evaluate generative fairness ‘$\mathcal{F}_{\tilde{x}_{T}}$’ based on the influence of tokens on textual guidance. Ideally, T2I model outputs should align with the input. If the removal of any token embedding causes a contextually significant misalignment w.r.t. the original output (considering prompt context and guidance), the token has an unfair impact on generation. An unreliable or intentionally-biased model may still generate diverse outputs of a given concept, depending on the learned relationships within conditional spaces. Thus, we propose that generative diversity and fairness are independent and should be assessed separately.
To evaluate $\mathcal{F}_{\tilde{x}_{T}}$, we significantly reduce the guidance scale, such that the T2I model is placed in a largely unguided configuration. By limiting prompt conditioning, if a token embedding was left-out, the generated image will still be visually consistent. A significant change in the output image under low guidance suggests that the left-out token has a large and unfair influence on the generated content. As visualized in Fig. 5, our $\mathcal{F}_{\tilde{x}_{T}}$ evaluations analyze the influence that tokens have on the generated image. In an intentionally-biased (unfair) model like [1], we see that removing the unreliable bias trigger token ‘ô’ causes a significant change, highlighting bias presence and provenance.
Given a token that causes unreliable model behavior $\tilde{x}_{T}\in\tilde{\mathbf{x}}$, we remove it from the input and generate images from $\mathcal{N}_{K}$ random noise samples to observe (on average) the influence of $\tilde{x}_{T}$ on semantic guidance w.r.t. the contextualized prompt $\tilde{\mathbf{x}}$. For the $k^{th}$ image, we can formalize generative fairness as
$\mathcal{F}_{\tilde{x}_{T_{k}}}=-\log(1-\frac{I_{\tilde{x}_{T_{k}}}\cdot I_{\tilde{\mathbf{x}}}}{I_{\tilde{x}_{T_{k}}}I_{\tilde{\mathbf{x}}}}),$ (7)
where $I_{\tilde{\mathbf{x}}}$ refers to the image generated using the prompt initially identified as causing unreliable model behavior. Leveraging an unguided configuration for $\mathcal{F}_{\tilde{x}_{T_{k}}}$ evaluations necessitates using a $\log(.)$ transformation to magnify smaller differences. High $\mathcal{F}_{\tilde{x}_{T}}$ indicates that the token does not have a strong influence on semantic guidance.
Figure 5: Generative Fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluation leaving out an unreliable token $\tilde{x}_{T}$ within the context of the unreliable prompt to assess its impact on generation. Each image represents one token removed from the prompt “a persôn drinking a coffee”. We see that while ‘ô’ persists in the input prompt, the output representation does not align with the input. When ‘ô’ is removed, the model behaves as expected. This demonstrates how sensitive triggers can have an unfair influence on guidance and thus, the generated image. Here, we use a high guidance example to illustrate the egregious impacts of intentional bias injections.
@@ 3.5 Intentional Bias Detection and Retrieval
We define $\mathcal{R}_{G}$, $\mathcal{R}_{L}$, $\mathcal{D}_{\tilde{x}_{T}}$ and $\mathcal{F}_{\tilde{x}_{T}}$ as general characteristics of T2I models. Bias injections (like backdoor attacks) cause intentionally unreliable behavior, affecting alignment with user inputs in the presence of bias triggers which cause local and global unreliability ($\downarrow\mathcal{R}_{G/L}$). Our fairness and diversity evaluations therefore enable effective intentional bias detection and retrieval of rare [1, 17] and natural language (NL) bias triggers [16], which signifies bias provenance. Figure 5 illustrates the behavior of a TPA-based, rare trigger [1], showing significant image changes when the trigger is removed. Figure 4 demonstrates how the BAGM [16] NL trigger “drink” yields uniform generated images when prompted. Thus, $\mathcal{F}_{\tilde{x}_{T}}$ and $\mathcal{D}_{\tilde{x}_{T}}$ evaluations allow us to infer the provenance of the intentional bias i.e., the trigger.
@@
Our reliability experiments offer initial insights into the presence/likelihood of intentional T2I model biases, acknowledging that ‘benign’ models may still show biased behavior for specific inputs, which limits binary detection approaches. Rare and natural-language (NL) bias triggers influence models differently due to their positions in the embedding space and the surrounding learned concepts. Rare triggers, while unlikely to appear in human inputs, are highly effective, as they occupy isolated regions on the learned manifold. In contrast, NL triggers are surrounded by similar concepts, making trigger retrieval increasingly difficult when similar terms are input. Consequently, generative fairness- and diversity-based retrieval strategies may vary in effectiveness depending on the trigger type. As shown in Fig. 4 and 5, we surmise that $\mathcal{D}_{\tilde{x}_{T}}$-based retrieval is more suited for NL triggers, while $\mathcal{F}_{\tilde{x}_{T}}$-based retrieval is more effective for rare triggers.
3.6 Implementation Details
[…]
Model Training dataset/s Learning
Rate
Text-Encoder Gen. Model Loss Function
Stable Diffusion V1.4 [60] laion2B-en
laion-high-resolution
laion-improved-aesthetics
laion-aesthetics v2 5+
0$\rightarrow$0.0001 CLIP ViT-L/14 2D cond. U-Net $\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(x_{t},t)^{2}_{2}]$
$\mathcal{L}_{DM}:=\mathbb{E}_{\epsilon(x),y,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))_{2}^{2}]$
Stable Diffusion V1.5
[58, 59]
laion2B-en
laion-high-resolution
laion-improved-aesthetics
laion-aesthetics v2 5+
0$\rightarrow$0.0001 CLIP ViT-L/14 2D cond. U-Net $\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(x_{t},t)^{2}_{2}]$
$\mathcal{L}_{DM}:=\mathbb{E}_{\epsilon(x),y,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))_{2}^{2}]$
Stable Diffusion V2.1 [61] laion2B-en
laion-high-resolution
laion-improved-aesthetics
laion-aesthetics v2 5+
0$\rightarrow$0.0001 OpenCLIP-ViT/H 2D cond. U-Net $\mathcal{L}_{DM}=\mathbb{E}_{x,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(x_{t},t)^{2}_{2}]$
$\mathcal{L}_{DM}:=\mathbb{E}_{\epsilon(x),y,\epsilon\sim\mathcal{N}(0,1),t}[\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))_{2}^{2}]$
BadT2I [17] *laion-aesthetics v2 5+ 0.00001 CLIP ViT-L/14 *2D cond. U-Net $\mathcal{L}_{Bkd-Obj}=\mathbb{E}_{\mathbf{z}_{b},\mathbf{c}_{b},\epsilon,t}[\epsilon_{\theta}(\mathbf{z}_{b,t},t,\mathbf{c}_{b\Rightarrow a,tr})-\hat{\epsilon}(\mathbf{z}_{b,t},t,\mathbf{c}_{b})_{2}^{2}]$
$\mathcal{L}_{Reg}=\mathbb{E}_{\mathbf{z}_{a},\mathbf{c}_{a},\epsilon,t}[\epsilon_{\theta}(\mathbf{z}_{a,t},t,\mathbf{c}_{a})-\hat{\epsilon}(\mathbf{z}_{a,t},t,\mathbf{c}_{a})_{2}^{2}]$
$\mathcal{L}=\lambda\cdot\mathcal{L}_{Bkd-Obj}+(1-\lambda)\cdot\mathcal{L}_{Reg}$
TPA [1] *laion-aesthetics v2 6.5+ 0.00001$\rightarrow$
0.0001
*CLIP ViT-L/14 2D cond. U-Net $\mathcal{L}_{BD}=\frac{1}{X}\underset{v\epsilon X}{\sum}d(E(y_{t}),\tilde{E}(v\oplus t))$
$\mathcal{L}_{Util.}=\frac{1}{X^{\prime}}\underset{w\epsilon X^{\prime}}{\sum}d(E(w),\tilde{E}(w))$
$\mathcal{L}=\mathcal{L}_{Util.}+\beta~{}\mathcal{L}_{BD}$
BAGM [16] *MF Dataset [62] 0.00001 *CLIP ViT-L/14 2D cond. U-Net $\mathcal{L}_{BD}=\frac{1}{N_{D}}\sum_{i=1}^{N_{D}}(y_{i}-\hat{y_{i}})^{2}$
Table 1: Implementation details and model information for the five unique models experimented with in this work. ‘*’ denotes intentionally-biased model datasets and target networks where applicable. All models use the the AdamW optimizer.
@@ 4 Results
Experimental Setup. We conduct our experiments in a grey-box setting, requiring access to internal model outputs (text-embeddings). Our evaluations do not require model weights or training information as is common in white-box settings [25]. To facilitate our evaluations, we deploy seven unique models, four of which are intentionally-biased/unreliable i.e.: (i, ii, iii) three off-the-shelf stable diffusion models (V1.4/1.5/2.1) [58, 60, 61, 34], (iv, v) two BAGM-based bias-injected models as per [16], (vi) a single-trigger, Target Prompt Attack (TPA) method based on [1] and, (vii) an object-manipulating BadT2I model based on [17]. Primary investigations leverage the Microsoft COCO dataset [63] as a source of input prompts. To model a different distribution of input prompts, we conduct an ablation study using prompts from the Google Conceptual Captions (GCC) dataset [64]. For primary image similarity calculations, we deploy a CLIP ViT-L/14 model to project images onto a common ViT feature space, calculating image similarity scores using these projections. Experiments with Google ViT (GViT) [65] and Facebook DeiT (FBViT) [66] image encoders are reported later. Additionally, we conduct an ablation study to identify any relationship between the number of generation steps and reported model reliability. As discussed prior, to support our diversity metric and the case of expected vs. unexpected lack of diversity, we present an ablation study to show shifts in diversity across levels of abstraction. To appropriately observe the behavior of intentionally-biased models, 10% of the test prompts contain a relevant input trigger depending on the model. For BAGM and TPA implementations, we train models based on provided code [16, 1]. The BadT2I model used for our evaluations is publicly available [17].
@@
Model Type $\varphi_{Mo}$ $\mathcal{P}_{G}(\varphi_{E}=\varphi_{Mo})$ $\varphi_{Mo}$ $\mathcal{P}_{L}(\varphi_{E}=\varphi_{Mo})$
@@
SD-V2.1 benign 0.1738 5.687 0.3983 2.742
BAGM NL-$\tilde{x}_{T}$ 0.1697 5.465 0.3755 2.236
Table 2: Global and local model reliability ‘$\mathcal{R}_{G/L}$$\varphi_{E}$-distribution comparison when generating images over $T=50$ denoising steps, using prompts from the MS COCO [63] dataset. ‘$\varphi_{Mo}$’ and ‘$\mathcal{P}_{G/L}(\varphi_{E}=\varphi_{Mo})$’ describe the Modal value and Mode, respectively.
@@
Global and Local Reliability. We attribute reliability to resistances to embedding perturbations ‘$\varphi_{E}$’. Across $\mathcal{R}_{G/L}$ experiments, when $\varphi_{E}$ reduces image similarity w.r.t. the original image below the threshold $\tau_{\varphi_{E}}\leq 0.9$, we record $\varphi_{s}$, the prompt and the token (for $\mathcal{R}_{L}$). To that end, we propose that the distribution function $\mathcal{P}_{G/L}(\varphi_{E})$ is viable for modeling both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$, reporting the Modal Value $\varphi_{Mo}$ and the Mode $\mathcal{P}_{G/L}(\varphi_{s}=\varphi_{Mo})$ in Table 2 and visualizing these distributions in Fig. 6.
As hypothesized, BadT2I and BAGM models are more sensitive to global $\varphi_{E}$, which can be seen through the shifting and squeezing of the distribution to the left, relative to the base model as reported in Table 2). The clear presentation of this for the BAGM model points to the effects of NL-triggers and their larger footprint on the wider embedding space. For the TPA, rare-trigger model, we see that it evades global reliability evaluations, which is testament to the globally stealthy nature of the proposed method [1]. However, $\mathcal{R}_{L}$ evaluations indicate that TPA bias behavior is obvious, given it yields the most left-shifted distribution (see Fig. 6 (b) and Table 2). This confirms the independence of our two reliability metrics and justifies the need for both $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations.
Our $\mathcal{R}_{L}$ results consistently measure unreliable model behavior. Unreliable models will be more sensitive to perturbed embeddings as highlighted by their left-shifted distributions and higher-magnitude peaks, which indicate a higher distribution of perturbation-sensitive tokens. Thus, our $\mathcal{R}_{G}$ and $\mathcal{R}_{L}$ evaluations can characterize extremely-biased model behavior and demonstrates an application of intentionally-biased model detection. Comparisons to baseline models demonstrates the effects that intentional bias injections can have on model reliability, based on the overall sensitivity to globally- and locally-applied $\varphi_{E}$. We view reliability from global and local perspectives to account for the notion that unreliable model behavior can manifest in different ways, through intentional bias injection methods or through unintentional biases that stem from training.
We propose evaluating T2I model reliability and use intentionally-biased models as benchmarks of known, unreliable/unfair cases. Applying similarity scores and thresholds is common in related backdoor detection works [56, 57]. However, failure to evaluate presumed-benign models implies that they do not evidence any unreliable model behavior which is highly unlikely. Exploiting local $\varphi_{E}$ and using only 10% triggered prompts, our $\mathcal{R}_{L}$ evaluation consistently identifies unreliable model behavior stemming from intentional bias injections. Unlike in our approach, [56, 57] only assume that a model is known to be biased (backdoored) and fail to evaluate on benign models. Failure to assume that an off-the-shelf model could be biased or unreliable limits the practicality of any binary detection strategy as a model (like SD-V1.4/1.5/2.1) may still operate with unintentional biases and act unreliably when exposed to certain input conditions. Therefore, we deemed it necessary to quantify reliability first.
@@
Fairness and Diversity. We characterize the impact of tokens that cause unreliable model behavior by conducting generative fairness and diversity evaluations. Unlike reliability evaluations that infer model behavior, this second evaluation stage quantifies acute model behavior when exposed to sensitive tokens. We present qualitative diversity evaluations for TPA and BAGM [1, 16] methods in Figs. 7(a) and 7(b). These qualitative findings support the initial motivations reported in Fig. 4. Figures 5 and 8 visualize our generative fairness results, using the TPA model [1], in which we exploit our ‘leave-one-out’ based approach. Conveniently, fairness and diversity evaluations are also viable in a black-box setup.
(a) SD-V1.5 vs. BAGM [16], input = “drink”
(b) SD-V1.5 vs. TPA [1], input = “ô”
Figure 7: Comparison of $\mathcal{D}_{\tilde{x}_{T}}$ evaluations for single token/concept inputs. For each sub-figure, the left hand-side shows generated images for the corresponding model. These images are then used to construct an image similarity matrix (right). We observe that in both BAGM [16] and TPA [1] examples, injecting bias into the model results in less diverse generated outputs. The significance of the reduction in diversity depends on the specificity of the concept used for bias injection. This is evidenced by the TPA [1] case which uses a rare trigger token.
Figure 8: Qualitative impact of the leave-one-out experiments that underpin generative fairness $\mathcal{F}_{\tilde{x}_{T}}$ evaluations. We highlight g=low vs. g=high guidance examples, which supports our decision to opt for a low-guidance scale in our evaluations. The top row of each example is the base model output. In both guidance scale setups, it is clear that removal of the bias trigger (highlighted red) causes the highest visual change in similarity w.r.t. the original output - the first column. However, when guidance is high, false-positives are more likely to appear, like when ‘peaceful’ is removed in the base model. This legislates using a low guidance for $\mathcal{F}_{\tilde{x}_{T}}$ evaluations. As shown in this case, true-positives demonstrate far more significant changes which is important for identifying bias provenance.
@@
We pose that the influence of a token on image guidance provides us with insights into rare trigger behavior - highlighting an example of this in Fig. 8. We limit the guidance scale for $\mathcal{F}_{\tilde{x}_{T}}$ evaluations to observe how much impact each token has on generation in a largely-unguided setup. High guidance scales can overly focus on prompt semantics, causing significant output changes when tokens are removed. This helps mitigate the impact of removing key objects from scenes e.g., if omitting the token ‘doctor’ from the prompt “a photo of a doctor.”
@@
Tokens with an unfair control over guidance will report significantly lower image similarities when left out, as evidenced in Table 3. The small variance in results is indicative of the low guidance scale, which necessitates using $\log(\cdot)$ in (7). Table 3 and Fig. 8 demonstrates the ability to retrieve the rare ‘ô’ trigger used for the TPA bias injection method [1]. Interestingly, while the diversity of the unreliable token ‘burger’ reported in the corresponding $\mathcal{D}_{\tilde{x}_{T}}$ column in Table 3 is significant, it also boasts an unfair influence over generation based on its position in the BAGM $\mathcal{F}_{\tilde{x}_{T}}$ column.
@@
These observations further evidence the influence that training has on learned representations. We demonstrate that through trigger retrieval, bias provenance can be established if a user suspects unreliable model behavior (quantified through $\mathcal{R}_{G/L}$ evaluations). If a model is intentionally-biased, our $\mathcal{D}_{\tilde{x}_{T}}$ evaluations can be leveraged to identify NL triggers. Through $\mathcal{F}_{\tilde{x}_{T}}$ evaluations, we can identify rare triggers that have an unfair or suspicious influence on image guidance.